Can someone help me on this problem?
I need to solve for x and check results:
9x + 7 = 5x -3
Thanks,
Sher

Answer by checkley71(8403)

9X+7=5X-3 OR 9X-5X=-3-7 OR 4X=-10 OR X=-10/4 OR X=-2.5
PROOF 9*-2.5+7=5*-2.5-3 OR -22.5+7=-12.5-3 OR -15.5=-15.5